# Linear combinations of latents in generative models: subspaces and beyond

### v5

@@ 1 Introduction
In this work, we demonstrate that successful generation from Gaussian latent vectors is strongly dependent on whether their broader statistical characteristics match those of Gaussian samples — a stronger condition than just having likely norms. We show that effective manipulation of latent spaces can be achieved by following the simple guiding principle of adhering to the modelling assumptions of the generation process. Our primary contributions are as follows:
  • We show that even if a latent leads to a valid object upon generation, it can fail to provide effective interpolation if it lacks the typical characteristics of samples — as diagnosed by statistical distribution tests.
@@ 3 Assessing validity of latents via distribution testing
Figure 4: Normality testing of latent vectors obtained from inversion. We show the LPIPS (Zhang et al., 2018) reconstruction errors (second row, in red) of 200 inverted randomly selected images across 50 random classes from ImageNet1k (Deng et al., 2009), the p-values of their inversions (third row), and rejection rates (bottom row) of the Kolmogorov-Smirnov normality test applied to the corresponding latent obtained from inversion under various step budgets. We use the diffusion model of Rombach et al. (2022), always using its maximum number of steps (999) for generation, and denote the 10th, 50th and 90th percentiles with black lines. The first row shows image reconstructions using its inversion at each budget, highlighted in red when the latent was rejected ($p<1e^{-3}$), with the interpretation that the characteristics of the latent were unlikely for a real sample according to the KS test. We note the strong correlation between inversion budgets providing low reconstruction errors and those for which the p-values of the latents are high — taking values likely to occur by chance for real samples. However, as we will see in Figure 5, there are still many latents with low reconstruction error yet extremely low p-value, and this often severely affects the quality of its interpolants.
@@ 3.1 The sample characteristics of seed latents affect their interpolations
Figure 5: Lack of normality is linked to failure of interpolants The left and middle panels shows LPIPS (Zhang et al., 2018) reconstruction error (after 999 generation steps) and the Kolmogorov-Smirnov p-value for all inversions presented in Figure 4, split into two plots due to the vast dynamic range of p-values. Although latents with high p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that also have low p-values. The right panel shows Q-Align visual quality scores (Wu et al., 2023) for spherical interpolants between pairs of inversions selected from matching ImageNet1k (Deng et al., 2009) image classes, demonstrating that choosing seed latents with both low reconstruction error ($<0.05$) and high p-values ($>1e^{-3}$) allows us to avoid low-quality interpolants that would arise when choosing seeds by reconstruction error alone. For reference, we include examples of interpolants at each visual quality level .
@@
Figure 4 demonstrates that high p-values are strongly correlated to latents obtained from inversions reproducing their original image with good quality. The p-value indicates the probability of observing the latent vector by chance given that it is drawn from the latent distribution $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. If this value is extremely low that is a strong indicator that the latent lacks characteristics expected of samples from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. In the figure we see that at low inversion budgets (with high reconstruction errors) most p-values are $1e^{-50}$ or lower, and then reach values you expect to see by chance for real samples (around $1e^{-3}$) at budgets where the reconstruction errors tend to be small. However, we will show that the p-value provides additional information about the quality of the latents than provided by reconstruction error alone.
We now examine how the p-values of two inversions relate to our ability to interpolate them with good quality. Figure 5 shows that although latents with high p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that have low p-values. Therefore, just because the process of going from object to latent and back reproduces the object, this does not necessarily mean that the resulting latent is well characterised as a sample from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. The lack of such characteristics can be inherited when manipulating the latent; in Appendix 8 we show that interpolants typically inherit low p-values from their seed latents. Figure 5 also demonstrates that if we beyond requiring inversions to reconstruct their original objects also require their p-values to be high — at levels expected of real samples — we are able to avoid many low quality interpolants downstream, by not using the rejected latents as seeds. In other words, the normality test helps us discover when the latents acquired from inversion lack characteristics of real samples, as this lack prevents them from being reliably used as seeds. Distribution testing the latents can be helpful for debugging the inversion setup — which typically involves solving an ODE — and assessing the settings used. In Appendix E we investigate the effectiveness of a range of other classic normality tests and tests based on the likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ as well as the likelihood of the norm statistic. We assess these methods across a suite of test cases including vectors with unlikely characteristics, and for which we know, through failed generation, violate the assumptions of the generative model.
@@ 4 Linear combinations of latents
  • Interpolation: $\bm{y}=v\bm{x}_{1}+(1-v)\bm{x}_{2}$, where $\bm{w}=[v,1-v]$ and $v\in[0,1]$.
  • Centroid Determination: $\bm{y}=\frac{1}{K}\sum_{k=1}^{K}\bm{x}_{k}=\sum_{k=1}^{K}\frac{1}{K}\bm{x}_{k}$, where $\bm{w}=[\frac{1}{K}]^{K}$.
  • Subspaces: Suppose we wish to build a navigable subspace spanned by linear combinations of $K$ latent variables. By performing the QR decomposition of $\bm{A}:=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{K}]\in\mathbb{R}^{D\times K}$ to produce a semi-othonormal matrix $\bm{U}\in\mathbb{R}^{D\times K}$ (as the Q-matrix), we can then define a subspace projection of any new $\bm{x}$ into the desired subspace via $s(\bm{x}):=\bm{U}\bm{U}^{T}\bm{x}=\bm{U}\bm{h}\in\mathbb{R}^{D}$. The weights $\bm{w}$ for a given point in the subspace $s(\bm{x})$ are given by $\bm{w}=\bm{A}_{\dagger}s(\bm{x})=\bm{A}_{\dagger}\bm{U}\bm{h}\in\mathbb{R}^{K}$ where $\bm{A}_{\dagger}$ is the Moore–Penrose inverse of $\bm{A}$. See the derivation of the weights and proof in Appendix D. One can directly pick coordinates $\bm{h}\in\mathbb{R}^{K}$, compute the subspace projection $\bm{y}=\bm{U}\bm{h}$, and then subsequently use the transport map (defined by Equation 3) to the latent space to obtain $\bm{z}$. In Figure 1 we use grids in $\mathbb{R}^{K}$ to set $\bm{h}$, used as above together with a basis (defined by $\bm{U}$) from a set of latents.
@@ 5.1 Interpolation and centroid determination
Method Accuracy FID Time
@@
Method Accuracy FID Time
@@ Acknowledgments
We thank Samuel Willis and Daattavya Aggarwal (University of Cambridge) for their valuable contributions. Samuel contributed with important ideas related to computational fluid dynamics simulation and the SLIDE model, and constructed the setup and plots used in Figure 7. Daattavya assisted in verifying proofs and offered comments that greatly improved the manuscript.

### v6

@@ 1 Introduction
In this work, we demonstrate that successful generation from latent vectors is strongly dependent on whether their broader statistical characteristics match those of samples — a stronger condition than just having e.g. likely norms. We show that effective manipulation of latent spaces can be achieved by following the simple guiding principle of adhering to the modelling assumptions of the generation process. Our primary contributions are as follows:
  • We show that even if a latent leads to a valid object upon generation, it can fail to provide effective interpolation if it lacks the typical characteristics of samples from the latent distribution — as diagnosed by statistical distribution tests.
@@ 3 Assessing validity of latents via distribution testing
Figure 4: Normality testing of latent vectors obtained from inversion. We show the LPIPS (Zhang et al., 2018) reconstruction errors (second row, in red) of 200 inverted randomly selected images across 50 random classes from ImageNet1k (Deng et al., 2009), the p-values of their inversions (third row), and rejection rates (bottom row) of the Kolmogorov-Smirnov normality test applied to the corresponding latent obtained from inversion under various step budgets. We use the diffusion model of Rombach et al. (2022), always using its maximum number of steps (999) for generation, and denote the 10th, 50th and 90th percentiles with black lines. The first row shows image reconstructions using its inversion at each budget, highlighted in red when the latent was rejected ($p<1e^{-3}$), with the interpretation that the characteristics of the latent were unlikely for a real sample according to the KS test. We note the strong correlation between inversion budgets providing low reconstruction errors and those for which the p-values of the latents are realistic — taking values likely to occur by chance for real samples. However, as we will see in Figure 5, there are still many latents with low reconstruction error yet extremely low p-value, and this often severely affects the quality of its interpolants.
@@ 3.1 The sample characteristics of seed latents affect their interpolations
Figure 5: Lack of normality is linked to failure of interpolants The left and middle panels shows LPIPS (Zhang et al., 2018) reconstruction error (after 999 generation steps) and the Kolmogorov-Smirnov p-value for all inversions presented in Figure 4, split into two plots due to the vast dynamic range of p-values. Although latents with high (realistic) p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that also have low p-values. The right panel shows Q-Align visual quality scores (Wu et al., 2023) for spherical interpolants between pairs of inversions selected from matching ImageNet1k (Deng et al., 2009) image classes, demonstrating that choosing seed latents with both low reconstruction error ($<0.05$) and high p-values ($>1e^{-3}$) allows us to avoid low-quality interpolants that would arise when choosing seeds by reconstruction error alone. For reference, we include examples of interpolants at each visual quality level.
@@
Figure 4 demonstrates that inversions having realistic p-values are strongly correlated with them reproducing their original image with good quality. The p-value indicates the probability of observing the latent vector by chance given that it is drawn from the latent distribution $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. If this value is extremely low that is a strong indicator that the latent lacks characteristics expected of samples from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. In the figure we see that at low inversion budgets (with high reconstruction errors) most p-values are $1e^{-50}$ or lower, and then reach values you expect to see by chance for real samples (around $1e^{-3}$) at budgets where the reconstruction errors tend to be small. However, we will show that the p-value provides additional information about the quality of the latents than provided by reconstruction error alone.
We now examine how the p-values of two inversions relate to our ability to interpolate them with good quality. Figure 5 shows that although latents with realistic p-values tend to have low reconstruction errors, there are many latents with low reconstruction errors that have low p-values. Therefore, just because the process of going from object to latent and back reproduces the object, this does not necessarily mean that the resulting latent is well characterised as a sample from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. The lack of such characteristics can be inherited when manipulating the latent; in Appendix 8 we show that interpolants typically inherit low p-values from their seed latents. Figure 5 also demonstrates that if we beyond requiring inversions to reconstruct their original objects also require their p-values to be realistic — at levels expected of real samples — we are able to avoid many low quality interpolants downstream, by not using the rejected latents as seeds. In other words, the normality test helps us discover when the latents acquired from inversion lack characteristics of real samples, as this lack prevents them from being reliably used as seeds. Distribution testing the latents can be helpful for debugging the inversion setup — which typically involves solving an ODE — and assessing the settings used. In Appendix E we investigate the effectiveness of a range of other classic normality tests and tests based on the likelihood $\mathcal{N}(\bm{x}_{\ast};\bm{\mu},\bm{\Sigma})$ as well as the likelihood of the norm statistic. We assess these methods across a suite of test cases including vectors with unlikely characteristics, and for which we know, through failed generation, violate the assumptions of the generative model.
@@ 4 Linear combinations of latents
  • Interpolation: $\bm{y}=\bm{w}^{T}\bm{X}$, where $\bm{X}=[\bm{x}_{1},\bm{x}_{2}]$, $\bm{w}=[w_{1},1-w_{1}]$ and $w_{1}\in[0,1]$.
  • Centroid Determination: $\bm{y}=\bm{w}^{T}\bm{X}$, where $\bm{X}=[\bm{x}_{1},\dots,\bm{x}_{K}],\bm{w}=[\frac{1}{K}]^{K}$.
  • Subspaces: Suppose we wish to build a navigable subspace spanned by linear combinations of $K$ latent variables. By performing the QR decomposition of $\bm{X}:=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{K}]\in\mathbb{R}^{D\times K}$ to produce a semi-orthonormal matrix $\bm{U}\in\mathbb{R}^{D\times K}$ (as the Q-matrix), we can then define a subspace projection of any new $\bm{x}$ into the desired subspace via $s(\bm{x}):=\bm{U}\bm{U}^{T}\bm{x}=\bm{U}\bm{h}\in\mathbb{R}^{D}$. The weights $\bm{w}$ for a given point in the subspace $s(\bm{x})$ are given by $\bm{w}=\bm{X}_{\dagger}s(\bm{x})=\bm{X}_{\dagger}\bm{U}\bm{h}\in\mathbb{R}^{K}$ where $\bm{X}_{\dagger}$ is the Moore–Penrose inverse of $\bm{X}$. See the derivation of the weights and proof in Appendix D. One can directly pick coordinates $\bm{h}\in\mathbb{R}^{K}$, compute the subspace projection $\bm{y}=\bm{U}\bm{h}$, and then subsequently use the transport map (defined by Equation 3) to the latent space to obtain $\bm{z}$. In Figure 1 we use grids in $\mathbb{R}^{K}$ to set $\bm{h}$, used as above together with a basis (defined by $\bm{U}$) from a set of latents.
@@ 5.1 Interpolation and centroid determination
Method Accuracy FID $\downarrow$ Time
@@
Method Accuracy FID $\downarrow$ Time
@@ Acknowledgments
We thank Samuel Willis and Daattavya Aggarwal (University of Cambridge) for their valuable contributions. Samuel contributed important ideas related to computational fluid dynamics simulation and the SLIDE model, and constructed the setup and plots used in Figure 7. Daattavya assisted in verifying proofs and offered comments that greatly improved the manuscript.